consecutive even integers: find four consecutive even integers whose sum is 340: consecutive even integers: find four consecutive even integers whose sum is 340
Answer by Alan3354(1916)

consecutive even integers: find four consecutive even integers whose sum is 340
----------------------
There are 2 ways to do this (maybe more, but 2 that I know)
The long way:
n + n+2 + n+4 + n+6 = 340
4n + 12 = 340
4n = 328
n = 82 (the 1st integer), 84, 86, 88
-------------
Quicker:
340/4 = 85
85 is the average, which is the value in the middle, so 2 will be less and 2 will be greater.
Less: 82, 84
Greater: 86, 88